OpenStudy (anonymous):

evaluate each polynomial for x= -1 and y=3 5x-y

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

@josemead @LoveYou*69 @Ozzy733 @Ozzy733 maybe you can help me I will give medal

OpenStudy (hyper):

At x = -1 and y = 3 , the polynomial equates to 5(-1) - 3 = -8

OpenStudy (anonymous):

thank you so much!

OpenStudy (hyper):

no worries any more questions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Evaluate each polynomial for x= -1 and y=3 6. x+3y 7. -7x+ x2y 8. 4y^2 + 11 x -16 please help me with this ones too!

OpenStudy (anonymous):

@Hyper please help!

OpenStudy (anonymous):

I will give you another medal

OpenStudy (hyper):

For 7), the polynomial equates to: -7(-1) + (-1)(2)(3) = 7 - 6 = 1 For 8), the polynomial equates to : 4(3 x 3) + 11(-1) -16 = 4(9) -27 = 27 - 27 = 0

OpenStudy (skullpatrol):

Do you understand what the phrase "evaluating a variable expression" means?

OpenStudy (skullpatrol):

evaluating a variable expression: Replacing each variable in the expression by a given value and simplifying the result.

OpenStudy (anonymous):

ooh okay

OpenStudy (skullpatrol):